OpenStudy (nuccioreggie):

Circle A has a diameter of 8 inches, a circumference of 25.12 inches, and an area of 50.24 square inches. The diameter of circle B is 3 inches, the circumference is 9.42 inches, and the area is 7.065 square inches. Part A: Using the formula for circumference, solve for the value of pi for each circle. (4 points) Part B: Use the formula for area and solve for the value of pi for each circle. (4 points) Part C: What observation can you make about the value of pi for circles A and B? (2 points)

OpenStudy (nuccioreggie):

Please help and. Ex ain the answer please

OpenStudy (nuccioreggie):

Explain the answer please

geerky42 (geerky42):

So you need to solve for \(\pi\). You can use these formulas: \(r = \dfrac{d}{2}\\~\\~\\ C = 2\pi r = \pi d\\~\\A = \pi r^2\)

geerky42 (geerky42):

Plug in values and solve for \(\pi\).

OpenStudy (nuccioreggie):

Formula for Area of Circle = (pi) * (radius)^2 Radius = (1/2) * (Diameter) Diameter of Circle A = 8 so Radius = (1/2) * 8 = 4 Area of Circle A = 50.24 (pi) = 3.14 Is that correct @geerky42

OpenStudy (nuccioreggie):

For A @geerky42

geerky42 (geerky42):

Well, pi does have to be somewhere close to \(3.14\)

OpenStudy (nuccioreggie):

Circumference = (pi)*(diameter) Circumference of Circle A = 25.12 and Diameter of Circle A = 8 so then it will be (pi) = (Circumference) divided by (Diameter) = 25.12 / 8 = 3.14 I got that for a and the another one is for B

OpenStudy (nuccioreggie):

@geerky42. Please come we have not finished my question and this is what i got so far m i right right Circumference = (pi)*diameter Circumference of Circle A = 25.12 and Diameter of Circle A = 8 so then it will be (pi) = Circumference divided by Diameter = 25.12 / 8 = 3.14 Formula for Area of Circle = (pi) * radius^2 Radius = (1/2) * Diameter Diameter of Circle A = 8 so Radius = 1/2 * 8 = 4 Area of Circle A = 50.24 (pi) = Area of Circle A divided by radius times radius= 3.14

OpenStudy (nuccioreggie):

@LegendarySadist

OpenStudy (nuccioreggie):

@amistre64

OpenStudy (nuccioreggie):

@pooja195

OpenStudy (nuccioreggie):

@texaschic101